OpenStudy (anonymous):

Find dy/dx. f(x)= √(5x^3-2x)

OpenStudy (anonymous):

\[f \left( x \right)=\left( 5x ^{3}-2x \right)^{\frac{ 1 }{ 2 }}\] now first diff. power only then inside the parentheses.

OpenStudy (anonymous):

I got that far, and I got to the next step, is this part correct?